We find the following information on NPNG (No-Pain-No-Gain) Inc. (18 marks total)

EBIT = $2,000,000

Depreciation = $250,000

Change in net working capital = $100,000

Net capital spending = $300,000

These numbers are projected to increase at the following super normal rates for the next three years, and 5% after the third year for the foreseeable future:

EBIT: 10%

Depreciation: 15%

Change in net working capital: 20%

Net capital spending: 15%

The firm’s tax rate is 35%, and it has 1,000,000 outstanding shares and $6,000,000 in debt. We have estimated the WACC to be 15%.

a.

Calculate the EBIT, Depreciation, Changes in NWC, and Net Capital Spending for the next four years. (8 marks)

b.

Calculate the CFA* for each of the next four years, using the following formula : CFA* = EBIT(1–T) + Depr – ΔNWC – NCS   (4 marks)

d.

Calculate the present value of growing perpetuity at Year 3.

(1 mark)

e.

Calculate the firm’s value at time 0 using the WACC of the firm as the discount rate. (Note that the first CFA* to be discounted is the cash flow from one year into the future.)

(3 marks)

f.

Calculate the firm’s equity value at time 0.

(1 mark)

g.

Calculate the firm’s share price at time 0
Answer #1

Answer:

A)

Sample Calculation

For year 2

EBIT=EBIT in previous year*(1+10%)=2000000*(1+10%)=$2200000

Depreciation =Depreciation in previous year*(1+15%)=250000*(1+15%)=$287500

NWC= NWC in previous year*(1+20%)=100000*(1+20%)=$120000

Net Capital = Net Capital in previous year*(1+15%)=300000*(1+15%)=$345000

Year 1 2 3 4
EBIT 2000000 2200000 2420000 2541000
Depreciation 250000 287500 330625 347156.25
NWC 100000 120000 144000 151200
Net Capital 300000 345000 396750 416587.5

B)

Year 1 2 3 4
EBIT 2000000 2200000 2420000 2541000
Depreciation 250000 287500 330625 347156.25
NWC 100000 120000 144000 151200
Net Capital 300000 345000 396750 416587.5
EBIT*(1-T) 1300000 1430000 1573000 1651650
CFA 1150000 1252500 1362875 1431018.75

D.

CF4=1431018.75

g=5%

r=15%

So PV at end of third year CFP3= CF1/(r-g)=1431018.75/(15%-5%)=$14310187.5

E.

Firm's PV at Year 0= CFA1/(1+15%)+CFA2/(1+15%)^2+CFA3/(1+15%)^3+CFP3/(1+15%)^3

Firm's PV at Year 0=1150000/(1+15%)+1252500/(1+15%)^2+1362875/(1+15%)^3+14310187.5/(1+15%)^3=$12252362.95

F.

Firm's Equity value at Time zero= PV of firm - Debt=12252362.95-6000000=$6252362.95

G. Firm share price= Firm's equity value / Outstanding shares=6252362.95/1000000=$6.25 per share